1) A scale is submerged at the bottom of a pool and then a 1kg block of aluminum swimming is placed on it. If the scale reads newtons, what does the scale read with the block on it? ON 2) A large rubber balloon has a mass of 0.3kg. To make it float in air, you begin inflating it with Helium. What volume will the balloon have when it barely floats in the air?
Answer #1

density of aluminium=2830 kg /m^3

volume of block = mass / density = 1 kg/ 2830 = 3.553 x 10^-4 m^3

Scale reading = Weight - buoyant force = mg - \rhoVg = 1 (9.8 ) - 1000 ( 3.553 x 10^-4) 9.8 = 9.8 - 3.48 = 6.32 N apprx

b) for floating , weight of ballon=  buoyant force

0.3 (9.8) = \rho Vg ( where \rho of air=1.2 kg/m^3)

V = 0.25 m^3